OpenStudy (anonymous):

find the y-intercept y=-1/6x-19/8

OpenStudy (anonymous):

at y intercept x=0 so put x=0 here y=-19/8 or compare it with y=mx+c form and you will gt c=-19/8:)

OpenStudy (anonymous):

@amandapollard28 medal??????????